# LATTE: Latent Trajectory Embedding for Diffusion-Generated Image Detection

### v1

@@ Abstract
The rapid advancement of diffusion-based image generators has made it increasingly difficult to distinguish generated from real images. This can erode trust in digital media, making it critical to develop generalizable detectors for generated images. Recent methods leverage diffusion denoising cues, but mainly focus on single-step reconstruction errors, ignoring the inherent sequential nature of the denoising process. In this work, we propose LATTE - Latent Trajectory Embedding - a novel approach that models the evolution of latent embeddings across several denoising timesteps. By modeling the trajectory of such embeddings rather than single-step errors, LATTE captures subtle, discriminative patterns that distinguish real from generated images. Each latent is refined by employing our latent-visual feature refinement module and aggregated into a unified representation. Afterwards, it is fused with the visual features and finally passed into a lightweight classifier. Our experiments demonstrate that LATTE surpasses the baselines on several established benchmarks, such as GenImage and DiffusionFake. Moreover, it demonstrates strong performance in cross-generator and cross-datasets settings, highlighting the potential of using the trajectory of latent embeddings for generated image detection. The code is available on the following link: https://github.com/AnaMVasilcoiu/LATTE-Diffusion-Detector.
@@ 1 Introduction
Diffusion-based generative models have fundamentally transformed the field of image generation [1, 2, 3, 4, 5, 6]. These models generate photorealistic content - such as portraits, landscapes, and complex scenes, by iteratively adding and then removing noise from data or latent representations, typically guided by a text prompt [7]. While this progress has unlocked transformative and creative applications, it has also facilitated the creation of fake images that are indistinguishable from authentic content. Such capabilities have already been exploited by malicious actors for various purposes, such as creating fraudulent impersonations of public figures [8, 9] or fabricating “evidence” in legal disputes [10, 11, 12], posing serious risks across political and social domains. These emerging threats underscore the urgent need for generated image detectors.
Recent efforts to detect diffusion-generated images [13, 14, 15, 16, 17, 18] have begun to exploit the distinctive signatures left by the generative process. For example, DIRE [13] hypothesizes that diffusion models can reconstruct synthetic images more accurately than real ones. Based on this insight, they proposed a novel representation that captures the error between an input image and its reconstruction, using this signal for classification. Another work, LaRE [14] proposes a more efficient alternative by performing only a single reverse diffusion step in latent space, rather than executing the full denoising trajectory, while still producing a discriminative error-based signal. Although these recent detection methods demonstrate strong performance, they overlook the inherent sequential nature of denoising. This limitation also hampers cross-generator generalization, since single-step errors and spatial artifacts fail to capture the consistent trajectory patterns shared across different generator pipelines.
In this paper, we introduce Latent Trajectory Embedding - LATTE, a novel framework that explicitly models the sequential evolution of latent representations across multiple denoising steps. During generation, diffusion models produce an image through a sequence of gradual denoising steps. At each step, the model applies a learned update that iteratively refines the sample, guiding it closer to the data manifold and producing an evolving representation. This iterative process defines a trajectory that reflects how the model interprets and refines the underlying content. We hypothesize that real images - whose details and textures can lie outside the model’s learned manifold - will often produce small inconsistencies between successive denoising steps, whereas fake images will follow smoother, more self-consistent trajectories. By sampling, refining, and aggregating these latent states, LATTE captures these irregularities that enable it to distinguish real from generated images.
@@
Specifically, given an image, we leverage a pretrained latent diffusion model to encode it into the latent space. Next, we apply forward noising and extract intermediate latent states during partial denoising at evenly spaced timesteps. We empirically find that evenly spacing the timesteps ensures coverage of early, mid, and late denoising stages, capturing the broad spectrum of the denoising dynamics. While this trajectory captures how the internal representation changes across steps, it does not indicate which image regions drive these changes. To enrich the LATTE representation, each latent is refined using a stack of transformer decoders that cross-attend to spatial features extracted from a pretrained image encoder. The refined sequence is then aggregated into a unified representation, fused with a global image feature, and passed to a lightweight classifier. This combination of latent dynamics and semantic cues enables LATTE to detect subtle inconsistencies indicative of generated content.
@@
We evaluate LATTE on two large-scale benchmarks, namely GenImage [19] and Diffusion Forensics [13], which include fake images generated by a diverse set of diffusion models. On average, our model surpasses current state-of-the-art methods, demonstrating strong detection performance. We report an average improvement of 3.1% on GenImage and a 7.1% in cross-domain settings on Diffusion Forensics. In particular, on one of the most challenging subsets of GenImage i.e., BigGAN [20] — the only GAN-based generator - LaTTE outperforms LaRE [14] by 13.3%, highlighting its cross-generator generalizability. In the cross-domain settings - for instance, the Bedroom partition of Dffusion Forensics - we observe a 16.3% gain over LaRE [14], underscoring LaTTE’s ability to generalize to specialized domains.
@@
In summary, our contributions are threefold: (1) We propose LATTE, the first diffusion-based feature that explicitly leverages the trajectory of latent states across multiple denoising steps. (2) We introduce a two-stage architecture that (i) samples and refines the latent trajectories via transformer decoders, and (ii) aggregates these refined embeddings into a compact, discriminative representation. (3) We demonstrate that LATTE achieves state-of-the-art performance and strong generalization across multiple image detection benchmarks, unseen generators, and domains.
@@ 2.1 Image Generation Models
Early methods for image synthesis were predominantly based on Generative Adversarial Networks (GANs) [21, 22, 20, 23, 24, 25], Variational Autoencoders (VAEs) [26, 27, 28, 29, 30, 31], and autoregressive models [32, 33, 34, 35], each introducing foundational ideas that shaped the evolution of generative modeling. GANs produce realistic images, but are hard to train and lack stable likelihood estimation. VAEs enable efficient inference and structured latent spaces but tend to generate blurry images. Autoregressive models offer precise likelihood modeling and detail but suffer from slow, sequential sampling, especially at high resolutions.
@@
To address earlier limitations, Denoising Diffusion Probabilistic Models (DDPMs) [1] introduced a generative process that reverses a gradual noising procedure, offering stable likelihood-based training and state-of-the-art image quality. Further advancements have explored improved sampling efficiency [2], accelerated solvers [36], architectural refinements [5, 4], and improved conditional generation with classifier-free guidance [37]. Latent Diffusion Models (LDMs) [3] improved scalability by operating in a compressed latent space learned via a variational autoencoder, enabling high-resolution synthesis at much lower cost. LDMs underpin popular models like Stable Diffusion, and have enabled extensions such as ControlNet [38] for spatial conditioning, SDXL [6] for ultra-high-resolution output, and LCM [39] for efficient few-step sampling.
In summary, diffusion-based models combine stable, likelihood-based training with fast, high-fidelity generation and have rapidly become the dominant paradigm in image generation. Their ability to produce highly realistic text-to-image content has significantly diminished the distinction between real and generated imagery, raising growing concerns around disinformation, media manipulation, and content authenticity. As a result, diffusion models now represent the primary focus of current research in generated image detection, as also addressed in this paper.
2.2 Generated Image Detection
@@
Early efforts in generated image detection primarily focused on GAN-generated content. Initial approaches relied on handcrafted features [40, 41], and later evolved into convolutional neural networks (CNNs) trained on large-scale datasets like FaceForensics++ [42]. Subsequent works explored more intrinsic traces of manipulation, including spectral artifacts in the frequency domain [43, 44] and inconsistencies in noise distributions [45, 46]. Although these approaches improved robustness across GANs, they have shown limited generalizability to diffusion-generated images.
@@
In response to the unique characteristics of diffusion-based generation, recent research has proposed detection strategies that leverage the inherent internal mechanics of the diffusion process. Some approaches focus on full image reconstruction: DIRE [13] introduced the idea of using DDIM [2] inversion error as a discriminative feature, while DRCT [16] uses a contrastive training objective on reconstructed images, and FIRE [47] emphasizes mid-frequency components in reconstruction errors. Other methods extract additional signals from the reverse denoising process. LaRE [14] improves efficiency by operating in latent space and using a single-step inversion. DNF [17] and SeDID [18] both leverage the predicted noise during denoising as an indicator of authenticity. On the contrary, in this work, we leverage the trajectory of latent states across multiple denoising steps as a more robust, discriminative image representation, aware of the evolution of the denoising process.
@@
Another line of research explores the use of powerful vision encoders, such as CLIP [48], which leverage vision-language alignment to capture rich semantic representations. These approaches use CLIP either as a frozen feature extractor with downstream classifiers [49], or in fine-tuned multi-modal frameworks that align image and text embeddings to capture inconsistencies in generated content [50, 51]. In this work, we also employ the vision encoder of CLIP in our experiments, alongside other large-scale vision encoders.
@@ 3 Methodology
In this section, we introduce our Latent Trajectory Embedding (LATTE) for diffusion-generated image detection. LATTE leverages the trajectory of the latent representations produced during the diffusion denoising process. In the next parts of this section, we first give a brief overview of the denoising process in latent diffusion models. Then, we continue by introducing LATTE and explaining how to extract and embed a sequence of latents into a unified trajectory representation. Finally, we explain how LATTE can be fused with visual features to enhance the detection of diffusion-generated images.
@@ 3.1 Preliminaries
Diffusion models define a Markov chain of diffusion steps that progressively add Gaussian noise to data until turning it into noise. In the literature, this is referred to as a forward noising process [1]. Specifically, starting from a clean image $x$, the forward chain gradually injects Gaussian noise over $T$ discrete steps:
@@
where $x_{t}$ is the noisy image at step $t$ and the schedule $\{\alpha_{t}\}$ controls the noise variance at each step. After $T$ steps, the image becomes nearly isotropic noise. In the reverse process, also defined as a Markov chain, the noisy image is gradually denoised to obtain the raw image. This process leverages a neural network $\epsilon_{\theta}(x_{t},t)$ to predict and remove this noise, defined as:
@@
To improve efficiency, latent diffusion models [3] first encode images into a lower-dimensional latent space via a pretrained VAE encoder $E_{\mathrm{VAE}}$, producing $z_{0}=E_{\mathrm{VAE}}(x_{0})$. The forward and reverse diffusion processes then operate on these latent codes $z_{t}\in\mathbb{R}^{d}$:
@@
$p_{\theta}(z_{t-1}\!\mid\!z_{t})$ $=\mathcal{N}\!\bigl{(}z_{t-1};\,\mu_{\theta}(z_{t},t),\,\Sigma_{\theta}(z_{t},t)\bigr{)},$ (4)
where $\epsilon_{\theta}$ is the neural network trained to predict the noise $\epsilon$. After denoising to $z_{0}$, a VAE decoder $D_{\mathrm{VAE}}$ reconstructs the final image $\hat{x}=D_{\mathrm{VAE}}(z_{0})$. Latent diffusion thus preserves high sample quality while significantly reducing computational and memory demands.
@@ 3.2 LATTE: Latent Trajectory Embedding
Diffusion models generate images by iteratively denoising a latent variable, gradually transforming noise into structure over a sequence of timesteps. To capture this progression, we propose LATTE, a representation that reflects how the latent evolves during the generative process. Rather than performing full reverse sampling, which is computationally expensive, we leverage the denoising model’s ability to operate at any time step $t$ and approximate the latent states with a single-step denoising per selected timestep.
Given an input image $x$, we first encode it into latent space using a pretrained VAE encoder: $z_{0}=E_{\mathrm{VAE}}(x)$. Next, for each selected timestep $t_{i}$, we simulate the forward diffusion process in one closed-form operation, producing a noisy latent:
@@
This one-step correction yields an estimate $\hat{z}_{t}$ that closely approximates the latent at $t$, while avoiding the overhead of a full reverse pass from $T$ to $t$. By repeating this forward-then-single-step reverse procedure for each of the $T$ timesteps $\{t,\ldots,T\}$ chosen to uniformly span the denoising schedule, we assemble the latent trajectory embeddings: $\mathcal{T}(x)=\{\hat{z}_{1},\hat{z}_{2},\dots,\hat{z}_{T}\}$ (see Figure 1). Crucially, this reduces latent feature extraction time by an order of magnitude compared to full reverse diffusion, as also discussed in [14].
@@ 3.3 Architecture Details
Our architecture combines two complementary feature streams — the LATTE sequence and visual semantics — to detect diffusion-generated images. As shown in Figure 2, it consists of two main stages: Latent–Visual Feature Refinement and Latent-Visual Fusion & Classification, which we describe next.
Figure 2: Overview of our proposed architecture using LATTE. It encompasses two main stages, namely, (1) Latent–Visual Feature Refinement where the LATTE is refined through stacks of $L$ cross-attention layers and (2) Latent-Visual Fusion & Classification for the final fusing with the visual features and predicting the output.
3.3.1 Latent–Visual Feature Refinement
To ground the denoising trajectory in the image content and highlight discriminative regions, each latent embedding is refined through cross-attention with spatial features extracted from a pretrained vision encoder.
Given an input image $x$, the encoder produces two outputs: (1) patch-level visual embeddings $\mathrm{V}\in\mathbb{R}^{N\times d}$, and (2) a global image token $\mathbf{v}_{\text{IMG}}\in\mathbb{R}^{d}$. The patch tokens $\mathrm{V}$ capture fine-grained spatial and semantic information from the image and are leveraged for the refinement of the LATTEs. The $\mathbf{v}_{\text{IMG}}$ token provides a holistic representation of the image and is used in the second stage.
Each latent embedding $\hat{z}_{t}$ in the trajectory $\mathcal{T}(x)$ is first flattened and linearly projected to match the dimensions $d$ of the visual features. The projected latents are then independently refined using a stack of $L$ transformer decoders, each consisting of a cross-attention layer followed by a feed-forward layer, with residual connections and layer normalization. Specifically, each latent $\hat{z}_{t}$ attends to the patch-level visual embeddings $\mathrm{V}$ using multi-head cross-attention (MHA) mechanism:
@@ 3.3.1 Latent–Visual Feature Refinement
where the keys and values $K,V\in\mathbb{R}^{N\times d}$ are both the visual features $\mathrm{V}$, $\mathbf{W}^{O}$ is the output projection layer, $h$ is the number of heads and $d$ is the dimension of the embeddings. Each $\hat{z}_{t}$ is refined through $L$ such layers, allowing it to align with different spatial features in the image independently of the other timesteps.
3.3.2 Latent-Visual Fusion & Classification
After refining each latent embedding through $L$ transformer decoder layers, we obtain a set of enriched embeddings ${\tilde{z}_{1},\ldots,\tilde{z}_{T}}$. To aggregate this sequence into a unified representation, we perform average pooling across all $T$ latents. Formally, we compute: $\tilde{z}_{\text{avg}}=\frac{1}{T}\sum_{i=1}^{T}\tilde{z}_{t}$.
This aggregated representation encodes how the latent embeddings transition through successive denoising steps, effectively summarizing the image’s reconstruction trajectory. Next, to incorporate a holistic semantic-level context, we concatenate $\tilde{z}_{\text{avg}}$ with the global image token $\mathbf{v}_{\text{IMG}}$:
$\mathbf{z}=[\tilde{z}_{\text{avg}}\mathbin{\|}\mathbf{v}_{\text{IMG}}]\in\mathbb{R}^{2d},$ (8)
Finally, we feed this joint embedding $\mathbf{z}$ into a lightweight linear classifier, which leverages this combined information to make a real-vs-generated prediction. By pooling refined latent embeddings and grounding them with the spatial semantics, our aggregation strategy effectively amplifies subtle artifacts that single-step or pixel-based methods tend to miss.

### v2

# Ana Vasilcoiu  1, Ivona Najdenkoska11footnotemark: 11,2, Zeno Geradts1,2, Marcel Worring1

@@ Abstract
The rapid advancement of diffusion-based image generators has made it increasingly difficult to distinguish generated from real images. This erodes trust in digital media, making it critical to develop generated image detectors that remain reliable across different generators. While recent approaches leverage diffusion denoising cues, they typically rely on single-step reconstruction errors and overlook the sequential nature of the denoising process. In this work, we propose LATTE - Latent Trajectory Embedding - a novel approach that models the evolution of latent embeddings across multiple denoising steps. Instead of treating each denoising step in isolation, LATTE captures the trajectory of these representations, revealing subtle and discriminative patterns that distinguish real from generated images. Experiments on several benchmarks, such as GenImage, Chameleon, and Diffusion Forensics, show that LATTE achieves superior performance, especially in challenging cross-generator and cross-dataset scenarios, highlighting the potential of latent trajectory modeling. The code is available on the following link: https://github.com/AnaMVasilcoiu/LATTE-Diffusion-Detector.
@@ 1 Introduction
Diffusion-based generative models have fundamentally transformed the field of image generation (Ho et al., 2020; Song et al., 2020; Rombach et al., 2022a; Nichol et al., 2021; Dhariwal & Nichol, 2021; Saharia et al., 2022; Podell et al., 2023; Midjourney, 2024; Black Forest Labs, 2025). These models generate photorealistic content - such as portraits, landscapes, and complex scenes - by iteratively adding and then removing noise from data or latent representations, typically guided by a text prompt (Rombach et al., 2022b). While this progress has unlocked transformative and creative applications, it has also facilitated the creation of fake images that are hard to visually distinguish from authentic content. Such capabilities have already been exploited by malicious actors, for instance, to create fraudulent impersonations of public figures (Twomey et al., 2023; de Rancourt-Raymond & Smaili, 2023) or fabricate “evidence” in legal disputes (Delfino, 2022; Sandoval et al., 2024; Koutras & Selvadurai, 2024). The challenge is also amplified by the growing landscape of image generation models, each introducing its own artifacts and characteristics. This underscores the urgent need for robust detectors able to distinguish real from generated images.
Recent efforts to detect generated images (Wang et al., 2023; Zhang & Xu, 2023; Ma et al., 2023; Luo et al., 2024; Ricker et al., 2024; Chen et al., 2024; Chu et al., 2024; Yan et al., 2025; Cheng et al., 2025) leverage distinctive signatures left by the generative process. Based on the hypothesis that diffusion models can reconstruct synthetic images more accurately than real ones, methods like DIRE (Wang et al., 2023) and LaRE (Luo et al., 2024) define novel representations that capture the error between an input image and its reconstruction. While achieving solid performance, these approaches rely on single-step representations and overlook the inherent sequential nature of denoising - a process that largely underlies the synthetic artifacts of fake images. We address this by treating the sequence of latent representations as a distinctive signature.
@@
In this paper, we introduce Latent Trajectory Embedding - LATTE, a novel approach that explicitly models the evolution of latent representations across multiple denoising steps. Namely, diffusion models generate images through a sequence of gradual denoising steps, where each learned update iteratively refines the sample toward the data manifold. This iterative process defines a trajectory that reflects how the model interprets and refines the underlying content. We hypothesize that real images, whose details and textures can lie outside the model’s learned manifold, will often produce small inconsistencies between successive denoising steps. On the contrary, fake images will follow smoother, more self-consistent trajectories aligned with the model’s generative prior. Specifically, given an image, we leverage a pretrained latent diffusion model to obtain its latent embedding. We apply standard forward noising and then extract intermediate latent states during the denoising at evenly spaced steps. This spacing provides a representative view of early, middle, and late denoising stages, capturing the full spectrum of the denoising dynamics. The resulting trajectory reflects how the internal representation evolves across steps, but it does not reveal which image regions drive these changes. To enrich the trajectory signal, we fuse each latent with visual features extracted from a pretrained image encoder using a stack of transformer decoders. The enriched sequence is subsequently aggregated into a compact representation, combined with global image features, and passed to a lightweight classifier. This combination of latent dynamics and semantic cues enables LATTE to leverage subtle inconsistencies indicative of generated content.
@@
We evaluate LATTE on well-established benchmarks for generated image detection, namely GenImage (Zhu et al., 2023b), Chameleon (Yan et al., 2025), and Diffusion Forensics (Wang et al., 2023). Our model surpasses current state-of-the-art methods, achieving an average improvement of 4.1% on GenImage over AIDE (Yan et al., 2025) and 7.1% in cross-domain settings on Diffusion Forensics over LaRE (Luo et al., 2024). In particular, on one of the most challenging subsets of GenImage i.e., BigGAN (Brock et al., 2018), LATTE outperforms the most competitive baseline by 9.5%, highlighting its cross-generator generalizability. In cross-domain settings - for instance, the Bedroom partition of Diffusion Forensics - we observe a 11.1% gain, underscoring LATTE’s robustness to specialized domains.
@@
In summary, our contributions are threefold: (1) We propose LATTE, the first diffusion-based embedding that explicitly leverages the trajectory of latent states across multiple denoising steps. (2) We introduce a two-stage architecture that (i) samples and enriches latent trajectories via transformer decoders and (ii) aggregates the latent embeddings into a compact and discriminative representation. (3) We demonstrate that LATTE achieves state-of-the-art performance and exhibits strong performance across diverse benchmarks, unseen generators, perturbations, and domains.
@@ 2 Related Work
Image Generation Models. Early methods for image generation were predominantly based on Generative Adversarial Networks (GANs) (Goodfellow et al., 2020; Karras et al., 2017; Brock et al., 2018; Choi et al., 2018; Park et al., 2019; Zhu et al., 2017), Variational Autoencoders (VAEs) (Kingma et al., 2013; Sohn et al., 2015; Zhao et al., 2017; Van Den Oord et al., 2017), and autoregressive models (Van den Oord et al., 2016; Parmar et al., 2018; Esser et al., 2021; Ramesh et al., 2021). GANs produce realistic images, but are hard to train and lack stable likelihood estimation. VAEs enable efficient inference and structured latent spaces but tend to generate blurry images. Autoregressive models offer precise likelihood modeling but suffer from slow, sequential sampling, especially at high resolutions.
@@
To address the limitations of early methods, Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al., 2020) introduced a generative process that reverses a gradual noising procedure, offering stable likelihood-based training and state-of-the-art image quality. Further advancements have explored improved sampling efficiency (Song et al., 2020), accelerated solvers (Karras et al., 2022), architectural refinements (Saharia et al., 2022; Nichol et al., 2021), and improved conditional generation with classifier-free guidance (Ho & Salimans, 2022). Latent Diffusion Models (LDMs) (Rombach et al., 2022a) improved scalability by operating in a compressed latent space learned via a variational autoencoder, enabling high-resolution generation at much lower cost. LDMs underpin popular models like Stable Diffusion, and have enabled extensions such as ControlNet (Zhang et al., 2023) for spatial conditioning, SDXL (Podell et al., 2023) for ultra-high-resolution output, and LCM (Luo et al., 2023) for efficient few-step sampling. Diffusion models now represent the primary focus of current research in generated image detection, as also addressed in this paper.
@@
Detection of Generated Images. Early efforts in generated image detection targeted GAN-generated content, starting with handcrafted features (Yang et al., 2019; Liy & InIctuOculi, 2018) and later advancing to convolutional neural networks (CNNs) trained on datasets like FaceForensics++ (Rossler et al., 2019). Subsequent works investigated intrinsic manipulation traces such as spectral artifacts in the frequency domain (Luo et al., 2021; Frank et al., 2020) and inconsistencies in noise distributions (Wang & Chow, 2023; Bai et al., 2024). While these approaches improved robustness across GANs, they demonstrated limited generalization to diffusion-generated images.
@@
To improve the generalizability of methods for detecting diffusion-based images, recent work has explored strategies that leverage the internal mechanics of the diffusion process. Some approaches focus on full image reconstruction: DIRE (Wang et al., 2023) introduced the idea of using DDIM (Song et al., 2020) inversion error as a discriminative feature, while DRCT (Chen et al., 2024) uses a contrastive training objective on reconstructed images. Other methods, like LaRE (Luo et al., 2024), improve efficiency by operating in latent space and using a single-step inversion. AIDE (Yan et al., 2025) incorporates low-level patch statistics and high-level semantics. In contrast, our method leverages the trajectory of latent states across denoising steps, capturing the evolution of the process as a more discriminative representation.
@@
Another line of research explores powerful vision encoders, such as CLIP (Radford et al., 2021), used either as a frozen feature extractor with downstream classifiers (Zhang et al., 2024), or in fine-tuned multi-modal frameworks aligning image and text embeddings to capture inconsistencies in generated content (Cozzolino et al., 2024; Li et al., 2024). We also employ CLIP’s vision encoder, alongside other large-scale vision encoders, to enrich our proposed latent trajectory embedding.
@@ 3 Methodology
In this section, we introduce our Latent Trajectory Embedding (LATTE) for generated image detection. First, we give a brief overview of the denoising process in latent diffusion models. Then, we continue by introducing LATTE and explaining how to extract and fuse a sequence of latents with visual features. Finally, we show how LATTE can be aggregated into a unified representation to enhance the detection of generated images.
@@ 3.1 Preliminaries
Diffusion models define a Markov chain of diffusion steps that progressively add Gaussian noise to data until turning it into noise. In the literature, this is referred to as a forward noising process (Ho et al., 2020). Specifically, starting from a clean image $x$, the forward chain gradually injects Gaussian noise over $T$ discrete steps:
@@
where $x_{t}$ is the noisy image at step $t$ and the schedule $\{\alpha_{t}\}$ controls the noise variance at each step. After $T$ steps, the image becomes nearly isotropic noise. In the reverse process, also defined as a Markov chain, the noisy image is gradually denoised to obtain the raw image. This backward chain leverages a neural network $\epsilon_{\theta}(x_{t},t)$ parameterized by $\theta$ to predict and remove this noise, defined as:
@@
To improve efficiency, latent diffusion models (Rombach et al., 2022a) first encode images into a lower-dimensional latent space via a pretrained VAE encoder $E_{\mathrm{VAE}}$, producing $z_{0}=E_{\mathrm{VAE}}(x_{0})$. The forward and reverse processes then operate on these latent codes $z_{t}\in\mathbb{R}^{d}$:
@@
$p_{\theta}(z_{t-1}\!\mid\!z_{t})$ $=\mathcal{N}\!\bigl(z_{t-1};\,\mu_{\theta}(z_{t},t),\,\Sigma_{\theta}(z_{t},t)\bigr).$ (4)
After denoising to $z_{0}$, a VAE decoder $D_{\mathrm{VAE}}$ reconstructs the final image $\hat{x}=D_{\mathrm{VAE}}(z_{0})$. Latent diffusion thus preserves high sample quality while reducing computational and memory demands.
@@ 3.2 LATTE: Latent Trajectory Embedding
In diffusion models, an image is reconstructed from noise by iteratively denoising latent variables over a sequence of timesteps (see Eqs. (3)–(4)). LATTE leverages the sequential structure of diffusion models by explicitly modeling how the latent embedding evolves across denoising steps. Instead of performing the full reverse chain, which is computationally expensive, we approximate intermediate states using single-step denoising at selected timesteps.
Given an input image $x$, we first encode it into latent space using a pretrained VAE encoder: $z_{0}=E_{\mathrm{VAE}}(x)$, as explained in section 3.1. Next, for each selected timestep $t$, we simulate the forward diffusion process in one closed‐form operation, producing a noisy latent:
@@
This one‐step correction yields an estimate $\hat{z}_{t}$ that closely approximates the latent at $t$, while avoiding the overhead of a full reverse pass from $T$ to $t$. By repeating this forward‐then‐single‐step reverse procedure for each of the $T$ timesteps $\{t,\ldots,T\}$ chosen to uniformly span the denoising schedule, we assemble the latent trajectory embeddings: $\mathcal{T}(x)=\{\hat{z}_{1},\hat{z}_{2},\dots,\hat{z}_{T}\}$, illustrated in Figure 1.
@@ 3.3 Architecture Details
Our architecture grounds the latent trajectory in visual context, ensuring that the latent representations remain tied to the image content. As illustrated in Figure 2, it unifies two complementary feature streams - the LATTE sequence and visual semantics - through two main stages: Latent–Visual Fusion and Latent-Visual Classifier.
Latent–Visual Fusion. Each latent embedding is enhanced through cross-attention with spatial features extracted from a pretrained vision encoder, to ground the denoising trajectory in the image content. Given an input image $x$, the vision encoder produces two outputs: (1) patch-level visual embeddings $\mathrm{V}\in\mathbb{R}^{N\times d}$, and (2) a global image token $\mathbf{v}_{\text{IMG}}\in\mathbb{R}^{d}$. The patch tokens $\mathrm{V}$ capture fine-grained spatial and semantic information from the image and are leveraged for the refinement of the LATTE representation. The $\mathbf{v}_{\text{IMG}}$ token provides a holistic representation of the image and is used in the second stage.
Each latent embedding $\hat{z}_{t}$ in the trajectory $\mathcal{T}(x)$ is first flattened and linearly projected to match the dimensions $d$ of the visual features $\mathrm{V}$. The projected latents are then independently enhanced using a stack of $L$ transformer decoders, each consisting of a cross-attention layer followed by a feed-forward layer, with residual connections and layer normalization. Specifically, each latent $\hat{z}_{t}$ attends to the patch-level visual embeddings $\mathrm{V}$ using multi-head cross-attention (MHA) mechanism:
@@
where the keys and values $K,V\in\mathbb{R}^{N\times d}$ are both the visual features $\mathrm{V}$, $\mathbf{W}^{O}$ is the output projection layer, $h$ is the number of heads and $d$ is the dimension of the embeddings. Each $\hat{z}_{t}$ is processed through $L$ such layers, allowing it to align with different spatial features in the image independently of the other timesteps.
Figure 2: Overview of our proposed architecture using LATTE. It encompasses two stages: (1) Latent–Visual Fusion, where the LATTE is fused with visual semantics through stacks of $L$ cross-attention layers, and (2) Latent-Visual Classifier for average aggregation and output prediction.
Latent-Visual Classifier. After enhancing each latent embedding through $L$ transformer decoder layers, we obtain a set of enriched embeddings $\tilde{z}_{1},\ldots,\tilde{z}_{T}$}. To aggregate this sequence into a unified representation, we perform average pooling across all $T$ latents: $\tilde{z}_{\text{agg}}=\frac{1}{T}\sum_{i=1}^{T}\tilde{z}_{t}$. Alternatively, we can perform $CLS$ token pooling where a special token ${z}_{\text{CLS}}$ is prepended to the sequence of latents ${z}_{\text{CLS}},\tilde{z}_{1},\ldots,\tilde{z}_{T}$}, processed with self-attention layers and then used as an aggregation $\tilde{z}_{\text{agg}}$. The aggregated representation encodes how the latent embeddings transition through successive denoising steps, effectively encoding the reconstruction trajectory. Next, to incorporate a holistic semantic-level context, we concatenate $\tilde{z}_{\text{agg}}$ with the global image token $\mathbf{v}_{\text{IMG}}$: $\mathbf{z}=[\tilde{z}_{\text{agg}}\mathbin{\|}\mathbf{v}_{\text{IMG}}]\in\mathbb{R}^{2d}$.
Finally, we feed this joint embedding $\mathbf{z}$ into a lightweight linear classifier, which leverages this combined information to make a real-vs-generated prediction. By pooling the latent embeddings and grounding them in image semantics, our aggregation strategy effectively amplifies subtle artifacts that single-step or pixel-based methods tend to overlook.